The test with 3 lengths of different Celani wires in a LENR-stick test cell is commencing in the water flow calorimeter. All the details are in the protocol document here: Protocol: Multi-Wire Test
The Experiment Log specifically for this test is linked from the document, along with the data field definitions.
It took us a bit longer to get this started this week because of some troubles with room temperature, air filters, and a leak around the pressure sensor. All that seems to be fixed and we are starting the loading phase, now. The loading will take a while since it requires manual intervention to go to each new level. The process was suggested by Celani.
- Loading at 5 bar
- Start at 75C, hold till resistance levels off
- Step up in 25C steps up to 450C, hold till resistance levels off (keep cell below hydrogen permeability of SS)
- Refresh pressure as needed due to absorption
- Cool down, drop pressure to 1 to 2 bar to allow flux out of wire on next heating cycle
- Heat up again to 350 or 400
- Watch for pressure rise indicating flux out of wire, and watch for excess heat.
Wish us luck and keep posting observations and suggestions.
The tests in the live data viewer are:
FC0404 Water Calorimeter
FC0405 LENR Stick: Multi-wire
This morning we are troubleshooting the water flow calorimeter to try to identify why the test cell appears to be running 10C higher than expected and the calorimeter shows almost a watt of excess. These two facts would be highly compelling on their own, but it appears that the exiting water temperature from the calorimeter was rising slowly over a couple of days before the hydrogen was even added. Our leading theory, at the moment, is that the cell temperature is due to changes in the foil wraps around it, and that thermistor 4 may be drifting- possibly due to moisture penetrating it.
To test this, we are turning off all the power to the cell, turning up the water flow to max, and watch to see if the sensor comes back in line with the others, or whether it retains a high bias. We'll keep you posted.
After doing the zero power test, we still saw Thermistor 4 reading higher than it should, so we decided to change the way things were working. We have seen thermistor drift on Thermistor-1 previously and just replaced it. This time we decided to re-engineer the way the sensors are installed. The old system had the thermistors sealed into heat shrink tubing with hot melt adhesive. The new system is tiny copper thermowells installed into the side of the silicone tubing. The thermistors are fixed intot he thermowells with a small amount of silicone adhesive along with a length of cotton sewing htread to act as a moisture wick to allow an exit path for any moisture finding its way near the sensor.
Below: Tiny copper thermowells made from .052" ID copper tube crimped and soldered on one end. They insert through the wall of the silicone tubing. They seal bey themselves, but we are adding a fine film of silicone seal to help make sure the seal is tight.
Below: A new thermowell inserted for the Thermistor 3 sensor.
Below: We see Thermistor 4's thermowell going into place. Beside it is the exposed trouble maker of the day wrapped in the heat shrink tubing. I took the sensor under a microscope, but didn't see any obvious signs of corrosion or any visible moisture with the plastic tubing.
Below: We see Thermistor 2 being installed with a dot of silicone and the cotton thread that will act as the wick.
Unfortunately, this means the test will be delayed till next week as the silicone adhesive cures over the weekend. We are hopeful that as long as both manners of installing the sensors provide accurate reading of the actual water temperature that it will not affect the calibration. That will have to be tested, though. We also, do not believe that the sensors were drifting during the calibration runs because of the tightness of the many different runs we did through the calibration steps. The data is available if anyone wants to validate that for us.
Think it will work?
We appear to have solved the drift problem with the new thermowells that are copper and much less water permeable. We re-zeroed the thermistors so they all related to each other nicely and had to debug a bit of debris in the flow regulating valve. Malachi did check the aluminum foild wrapping and found it to be loose and forming a spiral with multiple air layers that caused the internal temperatures to be higher for a given amount of power. While the insulation value is desirable, we decided to put it back the way it was during calibration. The next step is to check the calibration and make sure it reads appropriately. The challenge there is that we cannot know for sure that the wires are not giving off excess heat since they have now been exposed to Hydrogen, so if it appears to be off calibration at low powers, we will have to replace the cell with a control cell and re-calibrate.
Step 1: 3.5 watts input power and we will watch the power out and the internal temperature.
Last night we stepped the power up from 0.5 to 3.5 watts to try to gauge the slope of the calibrations and compare it to the original calibration. The result is that the slope is off a bit. That means we have to swap out the active LENR stick and replace it with a similarly configured one with a plain electric heater in it and run some calibration cycles on it.
Looks like next week before we get back to loading the active wires.